OpenStudy (anonymous):

Need help with 1 Question! (Question Inside) [MEDAL & FAN]

OpenStudy (anonymous):

@anwaarullah

OpenStudy (anwaarullah):

multiplyv2nd equation with 2vthen add both equations

OpenStudy (anonymous):

I know the final answer is (1,4) but I need the steps.

OpenStudy (anwaarullah):

3.25x-y=.75 multiply both sides with 2 6.50x-2y=3.75

OpenStudy (anwaarullah):

now get rid off y by adding two equations

OpenStudy (anwaarullah):

-4.5x-2y=-12.5 6.5x -2y =3.75

OpenStudy (anwaarullah):

add thses

OpenStudy (anonymous):

@mathmate

OpenStudy (anwaarullah):

2x= -8x x=4

OpenStudy (anwaarullah):

now put x=4in any equation to find y?

OpenStudy (anwaarullah):

-4.5x - 2y= -12.5 -4.5*4 -2y=-12.5

OpenStudy (mathmate):

@LetsLearn2000 If you would like to solve by substitution, as required by the question, this is what you do. 1. Make a choice of which variable out of x or y to be isolated and later substituted. Here, I choose y, because the coefficient of the second equation is -1. 2. Isolate y (i.e., y=....) from the second equation 3.25x-y=-0.75.............(2) transpose x to the right -y=-0.75-3.25x multiply throughout by -1 y=0.75+3.25x .............(2a) Substitute (2a) into the first equation where y is -4.5x-2y=-12.5.............(1) -4.5x-2(0.75+3.25x)=-12.5 Distribute -4.5x-1.5-6.5x=-12.5 simplify -11x=-12.5+1.5 -11x=-11 Divide by -11 x=1...............(1a) Put x into (2a) and solve for y y=0.75+3.25(x).........(2a) =0.75+3.25(1) =4 So the answer is (1,4) as you stated earlier.

OpenStudy (anonymous):

THANK YOU @mathmate

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (anonymous):

Thanks for the effort @anwaarullah